OpenStudy (anonymous):

write a quadratic equation in the varible x having the given number is type the equation in standard ax^2+bx+c=0 solution 6, only solution

OpenStudy (saifoo.khan):

x^2-12x+36

jimthompson5910 (jim_thompson5910):

x = 6 or x = 6 x-6 = 0 or x-6 = 0 (x-6)(x-6) = 0 x^2-12x+36 = 0

OpenStudy (anonymous):

(x-6)(x-6)=0 x^2-12x+36=0

OpenStudy (anonymous):

(x-6)^2 = x^2 -12 + 36